# The Lifespan of our Universe

### v1

@@ 1 Introduction
The age of our universe is of foundational importance in cosmology. Since the establishment of the Big Bang Theory, we know it is finite. In the $\Lambda$CDM model, together with the Cosmic microwave background (CMB) and Baryonic Acoustic Oscillation (BAO) and Supernova (SN) data, its age is now determined to be $13.8$ billion years [1]. The next question is about its future; more specifically, its lifespan. In this paper, we address this issue.
@@
If the observed dark energy is due to $\Lambda$ alone, earlier cosmological observations imply that $\Lambda>0$, so the universe continues to expand forever, with the equation of state (EoS) $w=-1$. However, the recent Dark Energy Survey (DES) and the Dark Energy Spectroscopic Instrument (DESI) measurements [2, 3] claim that, at 4.2 $\sigma$ level, the dark energy is dynamical; that is, the dark energy EoS $w\neq-1$. It is known that a simple ultra-light axion may explain this behavior [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. In addition to such an axion, the axion-dark energy (aDE) model [24] introduces a cosmological constant $\Lambda$  [^1] See also refs. [25, 26, 27, 28, 29, 30, 31]. Other implications on the $\Lambda<0$ case have been studied in refs. [32, 33, 34]., which happens to be a crucial feature. In applying this aDE model to the DES/DESI measurements, we find that the universe prefers to have a negative cosmological constant, i.e., $\Lambda<0$. In this case, $a$ reaches a maximum and then collapses to zero, ending in a big crunch. This big crunch defines the end of the universe.
In comparing to the DESI + CMB + BAO + SN data, the updated version of ref. [24] finds that there is a rather large degeneracy in the parameter space of the aDE model, where $\Lambda<0$ is most likely, implying that the universe has a finite lifespan. With the best-fit values for the parameters of the model, we find that the lifespan of the universe is 33.3 billion years; that is, the universe will end in about 20 billion years. It is encouraging that the lifespan of the universe can be quantitatively determined.
@@
In this paper, we briefly review the aDE model and how the DES/DESI data determines the parameters of the model. In particular, we discuss the likeliness that $\Lambda<0$. With the best-fit values of the model, we evolve the universe to the future and estimate the time when $a\to 0$, ending the universe in a big crunch. We discuss some issues and then conclude with some remarks.
@@ 2 The aDE Model
where $f$ is the so called decay constant and $\pi>\theta\geq 0$. If the universe starts at a random initial value $\theta=\theta_{i}\neq 0$, inflation will then lead to $\theta(\bm{x})=\theta_{i}=\phi_{i}/f$ everywhere. When the Hubble parameter $H\gg m_{\phi}$, the universe is essentially frozen at the misaligned initial state $\phi_{i}$ and $V(\theta_{i})$ contribute to the dark energy, i.e., the equation of state (EoS) $w=-1$. As $H\lesssim m_{\phi}$, $\phi$ starts to roll down along the potential towards $\phi=0$ and deposits the $\phi$ vacuum energy density into matter density (which drops like $a^{-3}$ where $a$ is the scale factor). The current dark energy is a combination of the cosmological constant $\Lambda$ and the axion field. The cosmological evolution is governed by Friedmann equations, which are
@@ 3 Best-fit values for the aDE model
There are 4 parameters in the model, namely $m_{\phi}$, $f$ or $\beta$, $\Lambda$ or $\Omega_{\Lambda}$, and the initial $\theta_{i}$ or $\phi_{i}$. The evolution of the axion and $w_{\rm DE}$ requires the knowledge of the (ordinary plus dark) matter density $\Omega_{\rm m}$ and the value of today’s Hubble parameter $H_{0}$. In the updated version of ref. [24], with $f=M_{\rm Pl}$, the DESI + CMB + BAO + SN data is used to constrain the 5 parameters ($m_{\phi},\phi_{i},\Omega_{m},\Omega_{\Lambda},H_{0}$), while $\Omega_{\phi}(m_{\phi},\phi_{i})$ (2.4) is determined.
Figure 1: The marginalized posterior distribution of $\Omega_{\Lambda}$ obtained by fitting the aDE model with a combination of BAO+SN+BBN+$\theta_{*}$+$t_{U}$ data, as described in ref. [24]. The blue curve is for the AdS $\Omega_{\Lambda}\leq 0$ case and the red curve is for the dS $\Omega_{\Lambda}>0$ case. The blue curve shows that the constraint on $\Omega_{\Lambda}$ is relatively unconstrained with the current DES/DESI data. As the blue ($\Omega_{\Lambda}\leq 0$) distribution is skewed towards higher $\Omega_{\Lambda}$, the corresponding best-fit value $\Omega_{\Lambda}=-1.61$ falls outside the range of the mean value $\langle\Omega_{\Lambda}\rangle=-0.69^{+0.70}_{-0.20}$.
The analysis aims to maximize the likelihood that observational data can be described by a background cosmology driven by the aDE model. Due to the limited available data, there is a relatively large degeneracy in the parameter space. In figure 1, only the posterior distribution of $\Omega_{\Lambda}$ is shown. It is convenient to separate the $\Omega_{\Lambda}>0$ side from the $\Omega_{\Lambda}\leq 0$ side. For $\Omega_{\Lambda}>0$, the best-fit value is consistent with $\Omega_{\Lambda}\lesssim 0.7$, implying that the axion field is not needed and $w=-1$ is acceptable. However, the matter density $\Omega_{m}\gtrsim 0.3$ seems to be on the high side.
Since the DESI meaurement claims $w\neq-1$ at 4.2 $\sigma$ level, let us focus the aDE model on the $\Omega_{\Lambda}\leq 0$ region. Here, the posterior distribution of $\Omega_{\Lambda}$ has a highly skewed degeneracy. As a result, the average value differs quite a bit from the best-fit value. In particular, as shown in ref. [20], setting $\Omega_{\Lambda}=0$ (i.e., the pure single axion case) is consistent with the data. However, it is much more likely that $\Omega_{\Lambda}$ is negative.  [^2] This point is also supported by a recent analysis ref. [35].
Here are the best-fit values for the aDE model obtained in ref. [24],
@@
Here $\Omega_{\phi,0}=\Omega_{\phi}(a=1)$ and $\Omega_{\rm DE,0}=\Omega_{\phi,0}+\Omega_{\Lambda}=0.72$. The EoS of dark energy $w_{\rm DE}$ may be obtained using (2.8), where we find that today $w_{0}=w_{\rm DE}(a=1)\simeq-0.738$. As mentioned in ref [24], this set of values provides a fit to the DES/DESI data that is comparable in quality to that of the $w_{0}w_{a}$ model. We shall use this set as a benchmark in calculating the lifespan of our universe. The consequences in using the mean values will be discussed later.
@@ 4.1 An approximate model
where $\Omega_{\phi,0}$ is a constant. Starting from $a=1$, we consider the effective $\Lambda$CDM model, where ${\bar{\Omega}}_{\rm m}$ now includes baryon matter, dark matter as well as the axion matter.
@@
and then approaches zero.
Note that in this approximation, we treat the axion as matter density for $t>t_{0}$ (before which it is dark energy) till the end (4.1). Therefore, the crunching phase is not simply time reversal of the expanding phase.
@@
Figure 2: The phase diagram for the Hubble parameter in the effective $\Lambda$CDM model. The arrow indicates the Hubble flow direction. The blue dot indicates a stable dS vacuum where the universe will expand exponentially with a constant $\mathcal{H}=\sqrt{\Omega}_{\Lambda}$. Here only matter and cosmological constant is included (4.2).
@@ 4.2 Estimating the lifespan
$\theta_{0}^{\prime}$ =±$=\pm\sqrt{\frac{3}{\beta^{2}}(1-\Omega_{\rm m})(1+w_{0})}\;,$ (4.6)
@@
This coupled system eq. (2.6) and eq. (4.7) can be solved numerically. Technically, to avoid the singularity, we set the threshold of the minimum scale factor as $a_{\rm min}=10^{-3}$, which terminates the calculation when $a<a_{\rm min}$ and set the lifespan $T\simeq t_{\rm max}$, where $a(t_{\rm max})\simeq a_{\rm min}$ [^4] Choosing a specific $a_{\rm min}$ is for the numerical purpose. Setting for example $a_{\rm min}=10^{-2}$ leads to numerical termination at the same $H_{0}t\simeq 2.29$, which makes no difference up to two decimal points. This is due to the rapid drop of $a$ towards the end, as shown in figure 4. Taking the best-fit parameter (3.1), we obtain the cosmic evolution as shown in figure 3. The scale factor reaches the maximum $a_{\rm max}\simeq 1.69$ at $H_{0}t\simeq 1.71$. Meanwhile, the Hubble parameter hits the zero at the same time and turn negative afterwards, which causes the crunch.
@@ 5 Discussion
  • As shown in Table 1 and Figure 4 in the extended version of ref. [24]. Because of the large degeneracy and the highly skewed probability distribution, the best-fit value $\Omega_{\Lambda}=-1.61$ is way outside the $1\sigma$ range of the mean value
@@
  • The Hubble tension is lurking at the background of our analysis as well as the DES/DESI analaysis [37]. If we bring in the axi-Higgs model [38, 39], $H_{0}$ can be uplifted by $2\%$$3\%$, relieving some of the Hubble tension. This will shorten the lifespan by a comparable amount.
@@ 6 Remarks
The best known case is the fuzzy dark matter (FDM) model, where an axion of mass around $10^{-22}~{}{\rm eV}$ is the source of the dark matter [49, 50]. To resolve some issues with the diversity of ultra faint dwarf galaxies, the introduction of a second axion of mass around $10^{-20}$ eV seems to be necessary [51]. Another even lighter axion of mass around $10^{-29}$ eV introduced in the so-called axi-Higgs model [38, 39] helps to explain the 7Li puzzle in BBN, the Hubble tension and the isotropic cosmic birefringence [52]. The aDE model introduces one with mass $10^{-33}$ eV. These 4 ultra-light axions have comparable decay constant $f\lesssim M_{\rm Pl}$. Presumably, they are Ramond-Ramond modes in string theory under the compactification to 3 spatial dimensions. If their existences are confirmed, these axions together with their properties will shed valuable insight into the string theory compactification.
Astrophysical/Cosmological Implications — In view of the dynamical dark energy impact on the evolution of the universe, it is interesting to re-analyze some of the dynamics involving astrophysical phenomena. As $a\to 0$, matter is being squeezed together. This is expected to enhance the formation of black holes, in particular the merging of black holes. Eventually, it is plausible that the universe ends in a single giant black hole. Including a growing radiation density as $a\to 0$ will make little difference.
Note that the scale factor $a$ has a periodic behavior and we consider only one cycle for our universe. As mentioned earlier, because of the nature of the axion field, the process is irreversible. Although very unlikely, one cannot rule out that there is a way for the universe to transition to the next cycle. However, this reincarnation will lead to a very different universe from our present one, not to a cyclic universe.
@@ 7 Summary
We start with $\Omega_{\rm m}+\Omega_{\phi,0}+\Omega_{\Lambda}=1$, where the matter density $\Omega_{\rm m}=0.28$ includes both the ordinary matter and the dark matter densities, while the dark energy density includes the axion energy density and the cosmological constant: $\Omega_{\phi,0}+\Omega_{\Lambda}=0.72$. Here, introducing a radiation density makes little difference. For the dark energy EoS $w=-1$, $\Omega_{\phi}=0$. For $w>-1$ and varying, the ultralight axion in the aDE model can explain its varying behavior, provided that $w\geq-1$. In fitting the DES/DESI data plus the CMB data, we determine the axion mass to be $m_{\phi}\simeq 2.93\times 10^{-33}$ eV and $\Omega_{\phi,0}\simeq 2.33$, so $\Omega_{\Lambda}\simeq-1.61$. This negative cosmological constant means that our universe contains the matter and the $\phi$ energy in an AdS space, resulting in its eventual contraction, ending in a big crunch. Defining the big crunch as the end of our universe allows us to estimate the lifespan of our universe to be about 33.3 billion years.
We emphasize that the determination of the lifespan of our universe depends on the recent observation that $w>-1$ and the validity of the aDE model. It is crucial that the DES/DESI observation is confirmed and the aDE model is rigorously tested. Fortunately, a number of projects measuring different aspects of the dark energy are forthcoming in the near future. We look forward to a more precise determination of the universe’s lifespan.

### v2

@@ 1 Introduction
[…]
@@
If the observed dark energy is due to $\Lambda$ alone, earlier cosmological observations imply that $\Lambda>0$, so the universe continues to expand forever, with the dark energy equation of state (EoS) w=1w=-1. However, the recent Dark Energy Survey (DES) and the Dark Energy Spectroscopic Instrument (DESI) measurements [2, 3] claim that, at 4.2 $\sigma$ level, the dark energy is dynamical; that is, the dark energy EoS $w\neq-1$. It is known that a simple ultra-light axion may explain this behavior [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. In addition to such an axion, the axion-dark energy (aDE) model [24] introduces a cosmological constant $\Lambda$  [^3] See also refs. [25, 26, 27, 28, 29, 30, 31]. Other implications on the $\Lambda<0$ case have been studied in refs. [32, 33, 34]., which happens to be a crucial feature. In applying this aDE model to the DES measurements, we find that the universe prefers to have a negative cosmological constant, i.e., $\Lambda<0$. In this case, $a$ reaches a maximum and then collapses to zero, ending in a big crunch. This big crunch defines the end of the universe.
In comparing to the BAO + SN data from DES and other constraints from CMB, BBN, age of the universe, ref.[24] finds that there is a rather large degeneracy in the parameter space of the aDE model, where $\Lambda<0$ is most likely, implying that the universe has a finite lifespan. With the best-fit values for the parameters of the model, we find that the lifespan of the universe is 33.3 billion years; that is, the universe will end in about 20 billion years. It is encouraging that the lifespan of the universe can be quantitatively determined.
@@
In this paper, we briefly review the aDE model and how the observational data determines the parameters of the model. In particular, we discuss the likeliness that $\Lambda<0$. With the best-fit values of the model, we evolve the universe to the future and estimate the time when $a\to 0$, ending the universe in a big crunch. We discuss some issues and then conclude with some remarks.
@@ 2 The aDE Model
where $f$ is the so called decay constant and $\pi>\theta\geq 0$. If the universe starts at a random initial value $\theta=\theta_{i}\neq 0$, inflation will then lead to $\theta(\bm{x})=\theta_{i}=\phi_{i}/f$ everywhere. When the Hubble parameter $H\gg m_{\phi}$, the universe is essentially frozen at the misaligned initial state $\phi_{i}$ and $V(\theta_{i})$ contribute to the dark energy, i.e., the equation of state (EoS) w=1w=-1. As $H\lesssim m_{\phi}$, $\phi$ starts to roll down along the potential towards $\phi=0$ and deposits the $\phi$ vacuum energy density into matter density (which drops like $a^{-3}$ where $a$ is the scale factor). The current dark energy is a combination of the cosmological constant $\Lambda$ and the axion field. The cosmological evolution is governed by Friedmann equations, which are
@@ 3 Best-fit values for the aDE model
There are 4 parameters in the model, namely $m_{\phi}$, $f$ or $\beta$, $\Lambda$ or $\Omega_{\Lambda}$, and the initial $\theta_{i}$ or $\phi_{i}$. The evolution of the axion and $w_{\rm DE}$ requires the knowledge of the (ordinary plus dark) matter density $\Omega_{\rm m}$ and the value of today’s Hubble parameter $H_{0}$. In the updated version of ref. [24], with $f=M_{\rm Pl}$, the BAO and SN data from DES [2] in combination with others are used to constrain the 5 parameters ($m_{\phi},\phi_{i},\Omega_{m},\Omega_{\Lambda},H_{0}$), while $\Omega_{\phi}(m_{\phi},\phi_{i})$ (2.4) is determined.
[…]
The analysis aims to maximize the likelihood that observational data can be described by a background cosmology driven by the aDE model. Due to the limited available data, there is a relatively large degeneracy in the parameter space. In figure 1 (left panel), only the posterior distribution of $\Omega_{\Lambda}$ is shown. It is convenient to separate the $\Omega_{\Lambda}>0$ side from the $\Omega_{\Lambda}\leq 0$ side  [^4] The distribution in the AdS case should be strictly cut off at $\Omega_{\Lambda}=0$, but the numerical package used (getdist) (mistakenly) extrapolates the edge to a slightly positive value to ensure normalizability. We have independently checked the MCMC simulations where $\Omega_{\Lambda}$ can run from AdS to dS. However, because this distribution is multi-modal with a strong attractor at $\Omega_{\Lambda}\sim 0.7$, these runs will end up getting stuck at the dS solution without exploring the AdS region, even though the latter actually yields a better fit (see Tab. I in ref. [24]). This is the reason we choose to separate these two scenarios.. For $\Omega_{\Lambda}>0$, the best-fit value is consistent with $\Omega_{\Lambda}\lesssim 0.7$, implying that the axion field is not needed and w=1w=-1 is acceptable. However, the matter density $\Omega_{m}\gtrsim 0.3$ seems to be on the high side.
Since the DES/DESI measurement claims $w\neq-1$ at 4.2 $\sigma$ level, let us focus the aDE model on the $\Omega_{\Lambda}\leq 0$ region. Here, the posterior distribution of $\Omega_{\Lambda}$ has a highly skewed degeneracy. As a result, the average value differs quite a bit from the best-fit value. In particular, as shown in ref. [20], setting $\Omega_{\Lambda}=0$ (i.e., the pure single axion case) is consistent with the data. However, it is much more likely that $\Omega_{\Lambda}$ is negative  [^5] A recent analysis ref. [36] independently argues that a negative $\Omega_{\Lambda}$ is most likely. (see Tab. I in [24] for a comparison of the best-fit values in the aDE model with $\Omega_{\Lambda}>0$ and $\Omega_{\Lambda}<0$).
[…]
@@
Here $\Omega_{\phi,0}=\Omega_{\phi}(a=1)$ and $\Omega_{\rm DE,0}=\Omega_{\phi,0}+\Omega_{\Lambda}=0.72$. The EoS of dark energy $w_{\rm DE}$ may be obtained using (2.8), where we find that today $w_{0}=w_{\rm DE}(a=1)\simeq-0.738$. As shown in figure 1, this set of values provides a fit to the DES data that is comparable in quality to that of the $w_{0}w_{a}$CDM model  [^6] This statement only applies for the aDE model with DES data as considered in ref. [24]. Other recent studies (e.g., [35, 39]) show that the “minimal thawing dark energy” model similar to our aDE model with $\Lambda=0$ is less favored than the $w_{0}w_{a}$CDM model when being constrained by DESI data.. For $a<0.23$, the aDE model yields w=1w=-1 while the $w_{0}w_{a}$CDM model predicts $w\lesssim-1.3$. This big difference will be tested in the near future. We shall use this set (3.1) as a benchmark in calculating the lifespan of our universe. The consequences in using the mean values will be discussed later.
@@ 4.1 An approximate model
[…]
@@
and then approaches zero. Note that in this approximation, we treat the axion as matter density for $t>t_{0}$ (before which it is dark energy) till the end (4.1). Therefore, the crunching phase is not simply time reversal of the expanding phase.
@@
[…]
@@ 4.2 Estimating the lifespan
$\theta_{0}^{\prime}$ $=\pm\sqrt{\frac{3}{\beta^{2}}(1-\Omega_{\rm m})(1+w_{0})}\;,$ (4.6)
@@
This coupled system eq. (2.6) and eq. (4.7) can be solved numerically. Technically, to avoid the singularity, we set the threshold of the minimum scale factor as $a_{\rm min}=10^{-3}$, which terminates the calculation when $a<a_{\rm min}$ and set the lifespan $T\simeq t_{\rm max}$, where $a(t_{\rm max})\simeq a_{\rm min}$ [^8] Choosing a specific $a_{\rm min}$ is for the numerical purpose. Setting for example $a_{\rm min}=10^{-2}$ leads to numerical termination at the same $H_{0}t\simeq 2.29$, which makes no difference up to two decimal points. This is due to the rapid drop of $a$ towards the end, as shown in figure 4. Taking the best-fit parameter (3.1), we obtain the cosmic evolution as shown in figure 3. The scale factor reaches the maximum $a_{\rm max}\simeq 1.69$ at $H_{0}t\simeq 1.71$. Meanwhile, the Hubble parameter hits the zero at the same time and turn negative afterwards, which causes the crunch. This happens roughly $11$ billion years from now.
@@ 5 Discussion
  • Note that the DESI (northern hemisphere) measurements are in good agreement with the DES (southern hemisphere) data set, so combining them yields the $w\neq-1$ result at $4.2\sigma$ significant level. As shown in figure 1, the best-fit dark blue curve (3.1) for the DES data set also fits well the DESI data set.
  • To gain a qualitative understanding on why a negative $\Omega_{\Lambda}$ is preferred, let us look at the redshift bin from DESI with the smallest error: $w\simeq-0.86$ at $a\simeq 0.77$ as shown in figure 1. The best fit red curve (i.e., with $\Omega_{\Lambda}>0$) is $1\sigma$ off from this data point. Here, $\Omega_{\phi,0}<0.72$ is limited. Qualitatively, a larger $\Omega_{\phi,0}$ tends to move $w$ closer to the central bin value. That is, a better fit (to the 2 lowest $z$ bins) requires a larger $\Omega_{\phi,0}$, which leads to a negative $\Omega_{\Lambda}$ and a big crunch.
  • As shown in Table 1 and figure 4 in the extended version of ref. [24], because of the large degeneracy and the highly skewed probability distribution, the best-fit value $\Omega_{\Lambda}=-1.61$ is way outside the $1\sigma$ range of the mean value
@@
  • The Hubble tension is lurking at the background of our analysis as well as the DES/DESI analaysis [41]. If we bring in the axi-Higgs model [42, 43], $H_{0}$ can be uplifted by $2\%$$3\%$, relieving some of the Hubble tension. Since the BAO and SN data are late time measurements, it may be more appropriate to use the late time Hubble value, i.e., $h=0.71$ instead of $h=0.67$. That is, $H_{0}$ shifts from $(14.5\,{\rm Gyr})^{-1}$ to $(13.8\,{\rm Gyr})^{-1}$, which will shorten the lifespan by about 1 billion years.
@@ 6 Remarks
The best known case is the fuzzy dark matter (FDM) model, where an axion of mass around $10^{-22}~{\rm eV}$ is the source of the dark matter [53, 54]. To resolve some issues with the diversity of ultra faint dwarf galaxies, the introduction of a second axion of mass around $10^{-20}$ eV seems to be necessary [55]. Another even lighter axion of mass around $10^{-29}$ eV introduced in the so-called axi-Higgs model [42, 43] helps to explain the 7Li puzzle in BBN, the Hubble tension  [^9] See also ref. [56] for another possibility. and the isotropic cosmic birefringence [57, 58, 31]. The aDE model introduces one with mass $10^{-33}$ eV. These 4 ultra-light axions have comparable decay constant $f\lesssim M_{\rm Pl}$. Presumably, they are Ramond-Ramond modes in string theory under the compactification to 3 spatial dimensions. If their existences are confirmed, these axions together with their properties will shed valuable insight into the string theory compactification.
Astrophysical/Cosmological Implications — In view of the dynamical dark energy impact on the evolution of the universe, it is interesting to re-analyze some of the dynamics involving astrophysical phenomena. As $a\to 0$, matter is being squeezed together. This is expected to enhance the formation of black holes, in particular the merging of black holes. Eventually, it is plausible that the universe ends in giant black holes. Including a growing radiation density as $a\to 0$ will make little difference.
Note that the scale factor $a$ has a periodic behavior and we consider only one cycle for our universe. As mentioned earlier, because of the nature of the axion field, the process is irreversible. Although very unlikely, one cannot rule out that, in the presence of quantum effects, there is a way for the universe to transition to the next cycle. However, this reincarnation will lead to a very different universe from our present one, not to a cyclic universe.
There are 3 types of singularities we encounter in the study of gravity : black holes, the origin of the universe and the crunch at the end of the universe. As the universe is collapsing, one can imagine that matters are push together to form a giant black hole, which in turn shields/hides the crunch singularity.
@@ 7 Summary
We study a very simple axion plus $\Lambda$ model with $\Omega_{\rm m}+\Omega_{\phi,0}+\Omega_{\Lambda}=1$, where the matter density $\Omega_{\rm m}=0.28$ includes both the ordinary matter and the dark matter densities, while the dark energy density includes the axion energy density and the cosmological constant: $\Omega_{\phi,0}+\Omega_{\Lambda}=0.72$. Here, introducing a radiation density makes little difference. For the dark energy EoS w=1w=-1, $\Omega_{\phi}=0$. For w>1w>-1 and varying, the ultralight axion in the aDE model can explain its varying behavior, provided that $w\geq-1$. In fitting the DES data in combination with other constraints from CMB, BBN and age of the universe, we determine the axion mass to be $m_{\phi}\simeq 2.93\times 10^{-33}$ eV and $\Omega_{\phi,0}\simeq 2.33$, so $\Omega_{\Lambda}\simeq-1.61$. This negative cosmological constant means that our universe contains the matter and the $\phi$ energy in an AdS space, resulting in its eventual contraction, ending in a big crunch. Defining the big crunch as the end of our universe allows us to estimate the lifespan of our universe to be about 33.3 billion years.
We emphasize that the determination of the lifespan of our universe depends on the recent observation that w>1w>-1 at small redshift and the validity of the aDE model. It is crucial that the DES/DESI observation is confirmed and the aDE model is rigorously tested. Fortunately, a number of projects measuring different aspects of the dark energy are forthcoming in the near future. We look forward to a more precise determination of the universe’s lifespan.